there are 120 members in a club
there are 1.5 more girls than boys in this club
how many boys are there

Answer by Alan3354(69443)

there are 120 members in a club
there are 1.5 more girls than boys in this club
how many boys are there
----------------
b + g = 120
g = b + 1.5
----
b + (b+1.5) = 120
2b + 1.5 = 120
2b = 118.5
b = 59.25 boys
g = 60.75 girls
---------
I don't know how that will work.